Five years ago you purchased a $1,000 U.S. Treasury bond for $920. At the time of...

  1. Five years ago you purchased a $1,000 U.S. Treasury bond for $920. At the time of purchase it had a 12- year remaining maturity and a 6% coupon (paid semiannually). Market interest rates have decreased 2% since then. What is the current market value of your bond?
Answer #1

Hello Sir/ Mam

YOUR REQUIRED MARKET VALUE IS $1,058.45.

Five years ago:

PV = $920

FV = $1000

n = 24

PMT = $60/2 = $30

Hence, we use excel to calculate YTM using "=RATE(24,30,-920,1000,0)" and it comes out to be approximately 3.5%. Hence, annual rate = 7%

Now,

YTM = (7% - 2%)/2 = 2.5%

FV = $1000

n = 24 - 10 = 14

PMT = $60/2 = $30

Now, using excel formula, "=PV(2.5%,14,-30,-1000,0)", we get the answer that the present market price equals $1,058.45.

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.